OpenStudy (anonymous):

How do you solve 2(x-3)^2=98 algebraically? My answer key says the answer is 10, but I think it is wrong

OpenStudy (anonymous):

thats right its 10

OpenStudy (anonymous):

how do you get that

OpenStudy (anonymous):

(x-3)^2 = 49 (x-3) = +-7 x = 10 or -4

OpenStudy (anonymous):

no its =98